# Invariant derivations and trace bounds

### v3

@@ 11. Type $C$
Note that the eigenvalues of an element of $\operatorname{Sp}_{2n}$ coming from the compact form have $\bar{s}_{j}=1/s_{j}$, so each $t_{j}$ is real. It follows that the power sums $p_{d}=\sum_{i}t_{i}^{d}$ have $p_{1}^{2}$ and $p_{2}$ positive. The expression $\sum t_{j}^{2}+\sum t_{j}t_{k}$, which is the degree 2 complete homogeneous symmetric function in the $t_{j}$’s, so it is $(p_{1}^{2}+p_{2})/2\geq 0$. Then (11.1) gives the required bound $\operatorname{Tr}\operatorname{Ad}(g)\geq-(\operatorname{rank}G)$, proving Theorem 2.3 for type $C$.
@@
The claim for type $C_{n}$ in Proposition 2.5 now follows directly from the $D_{n}$ case of Theorem 2.3.
@@ 12. Type $B$
The highest short root representation $\rho$ of $\operatorname{SO}(2n+1)$ is the natural representation $k^{2n+1}$, so
@@ 13. Components other than the identity
To complete the proof of Theorem 2.3, it remains to compute $\underline{m}(s)$ and $\overline{m}(s)$ when $s\neq 1$.
In the following results, we consider a compact Lie group $G$ that is not connected. We leverage the following slight sharpening of an old result of de Siebenthal [dS56], the elements of which already appear in [Rai97, §2].
@@
Let $C$ be a connected component of a compact Lie group $G$. There exists a $c\in C$ of finite order and a torus $T$ of $Z_{G}(c)$, such that $C$ is a union of $G^{\circ}$-conjugates of $cT$.
Proof.
The theorem on page 57 of [dS56] says that for any element $x\in C$ and every maximal torus $T$ of $Z_{G}(x)$, $C$ is a union of $G^{\circ}$-conjugates of $xT$. Because $Z_{G}(x)\subset G$ is closed, it is compact, so the quotient $Z_{G}(x)/Z_{G}(x)^{\circ}$ is finite, and thus there is some power $m$ of $x$ that lies in $Z_{G}(x)^{\circ}$. Since $x^{m}$ is central in $Z_{G}(x)^{\circ}$, we find that $x^{m}$ lies in every maximal torus of $Z_{G}(x)$ and thus in $T$. Since $T$ is divisible, there is a $t\in T(\mathbb{R})$ such that $t^{-m}=x^{m}$. We take $c:=tx$, which has order $m$. By de Siebenthal, $C$ is a union of $G^{\circ}$-conjugates of $xT=cT$. Finally, note that $T$ is contained in $Z_{G}(x)\cap Z_{G}(t)\subseteq Z_{G}(c)$. ∎
Given a connected component $C$, we write $c$ for the element provided by the lemma and $T_{c}$ for a maximal torus in $Z_{G}(c)$ containing the $T$ provided by the lemma.
Lemma 13.2.
If $G^{\circ}$ is not a torus, then $\dim(T_{c})>0$.
Proof.
It is certainly the case that $c$ normalizes some maximal torus $T$ of $G^{\circ}$, and thus normalizes the Weyl group of $T$. In particular, we may multiply it by some element of $N_{G^{\circ}}(T)$ so that it acts as a diagram automorphism on $T$. Since this acts as a permutation on a suitable basis of the character lattice of $T/Z(G^{\circ})$, it centralizes a nontrivial torus in $T$ as required. ∎
Lemma 13.3.
The centralizer of $T_{c}$ in $G^{\circ}$ is a (maximal) torus.
Proof.
Suppose otherwise. Then $c$ acts on the centralizer $Z_{c}$ of $T_{c}$ and thus on the quotient $Z_{c}/T_{c}$. Applying the previous lemma to $Z_{c}/T_{c}\rtimes\langle cT_{c}\rangle$ gives a torus $T^{\prime}_{c}\in Z_{c}/T_{c}$ which is centralized by $c$. Since some power of $c$ is in $T_{c}$, the action on $Z_{c}/T_{c}$ has finite order, and thus so is the action on the preimage of $T^{\prime}_{c}$ in $Z_{c}$. By construction, the action on the preimage is unipotent on the character group, and thus must be trivial, contradicting maximality of $T_{c}$. ∎
Note that since by construction $T$ centralizes $T_{c}$, any element of $cT$ centralizes $T_{c}$ and thus has the same associated torus.
Extremizing class functions
One approach to extremizing class functions on nonidentity components, which we do not follow in this paper, would be to mimic our approach for connected groups. The quotient scheme $\operatorname{Class}_{G^{\circ}}(C)$ may be identified with the quotient $cT_{c}/N_{G^{\circ}}(cT_{c})$, in which the identity component $T_{c}$ of $N_{G^{\circ}}(cT_{c})$ acts trivially. This generalized Weyl group $W_{c}$ is now acting on a torsor over a torus rather than a torus itself, but this was not essential in proving that critical points can be detected using invariant derivations. The construction of an invariant operator $\mathcal{D}_{A}$ applies mutatis mutandis, and thus if $cT_{c}/W_{c}$ can be identified with an open subset of affine space, we can again construct a basis of invariant derivations. Similarly, the action of $\pi_{1}(cT_{c}).W_{c}$ on the universal cover is by Euclidean transformations, so the results of section 6 apply directly, again assuming that the quotient is smooth. This holds up to isogeny: if $G^{\circ}$ is simply connected, then $\pi_{1}(cT_{c}).W_{c}$ is an affine Weyl group [Mat02, (5.2)], and thus $cT_{c}.W_{c}$ is an affine space. (In fact, this tells us that it may be identified with $\operatorname{Class}(\hat{G}_{c})$ for some simply connected group $\hat{G}_{c}$; a case by case analysis shows that when $G^{\circ}$ is simply laced and $c$ is as above, $\hat{G}_{c}$ is the Langlands dual of $(G^{\circ})^{c}$.)
In the case of the adjoint representation, however, there is a simpler (if kludgier) approach that does not require us to develop the full theory of disconnected groups (e.g., figuring out how to express characters of $\langle C\rangle$ in terms of the coordinates on $cT_{c}/W_{c}$). In fact, it turns out that we can reduce to the adjoint representation of a suitable (connected!) subquotient of $(G^{\circ})^{c}$, and thus to a calculation we have already done.
A subquotient $H$
Lemma 13.4.
Let $G$ be a compact Lie group with maximal torus $T$, let $w\in N_{G}(T)$, and let $\rho:G\to\operatorname{GL}(V)$ be a representation of $G$. Then for all $t\in T$,
$\operatorname{Tr}(\rho(wt))=\operatorname{Tr}(\rho(wt)|_{V^{[w,T]}}.$
@@
Over $\mathbb{C}$, $V$ splits as a direct sum of weight spaces $V_{\lambda}$. If $w\cdot\lambda\neq\lambda$, then the action of $w$ permutes the corresponding weight spaces, so those summands together contribute 0 to the trace. We thus find that
$\operatorname{Tr}(\rho(w))=\sum_{\lambda:w\cdot\lambda=\lambda}\operatorname{Tr}(\rho(w)|_{V_{\lambda}})=\operatorname{Tr}(\rho(w)|_{W}),$
where $W$ is the direct sum of the weight spaces corresponding to $w$-invariant weights. A weight is $w$-invariant if and only if $\lambda(wtw^{-1})=\lambda(t)$ for all $t\in T$, if and only if $\lambda(wtw^{-1}t^{-1})=1$ for all $t\in T$, if and only if $\lambda|_{[w,T]}=1$. In other words, $W$ is nothing other than the subspace of $V$ invariant under $[w,T]$ as required. ∎
In particular, when optimizing a character over a coset $wT$, we may restrict to the subspace $V^{[wT,T]}$. This is naturally a representation of a subquotient of $G$, namely $N_{G}([wT,T])/[wT,T]$. Call the identity component of this group $H$. Note that $wT\subset N_{G}([wT,T])/[wT,T]$, so any element of $wT$ normalizes $H$. Also, the Lie algebra of $H$ is nothing other than the $[wT,T]$-invariant subspace of $\mathfrak{g}$, so by the same discussion, its roots (i.e., the nontrivial weights of $\mathfrak{g}^{[wT,T]}$) are nothing other than the $w$-invariant roots of $G$, while its maximal torus is $H_{0}(\langle w\rangle;T)$.
@@
There is an element $w_{0}\in wT$ that centralizes $H$.
@@
The quotient $T/[wT,T]=H_{0}(\langle w\rangle;T)$ is a maximal torus of $H$, and any element of $wT$ fixes that torus pointwise. It follows that the image of $wT$ in $\operatorname{Aut}(H)$ is contained in the centralizer of $T/[wT,T]$, so is contained in the corresponding maximal torus of $\operatorname{Ad}(H)$. The map from $T$ to that maximal torus is surjective, and thus the preimage of the identity is nonempty, so in particular contains an element $w_{0}$ as required. ∎
This leads to a particularly nice conclusion when $V$ is the adjoint representation.
@@
$\operatorname{Tr}(\operatorname{Ad}_{G}(wt))$ $=\operatorname{Tr}(\operatorname{Ad}_{H}(t))+\operatorname{Tr}(\operatorname{Ad}_{G}(w)|_{\operatorname{Lie}([wT,T])})$
$=\operatorname{Tr}(\operatorname{Ad}_{\operatorname{Ad}(H)}(t))+\operatorname{Tr}(\operatorname{Ad}_{G}(w)|_{\mathfrak{t}})-\operatorname{rank}(\operatorname{Ad}(H)).$
@@
Let $H^{+}$ be the normalizer of $[wT,T]$, so that its Lie algebra is nothing other than $\mathfrak{g}^{[wT,T]}$, and thus
@@
$\operatorname{Tr}(\operatorname{Ad}_{H^{+}}(wt))$ $=\operatorname{Tr}(\operatorname{Ad}_{H}(wt))+\operatorname{Tr}(\operatorname{Ad}_{H^{+}}(wt)|_{\operatorname{Lie}([wT,T])})$
$=\operatorname{Tr}(\operatorname{Ad}_{H}(wt))+\operatorname{Tr}(\operatorname{Ad}_{G}(wt)|_{\operatorname{Lie}([wT,T])}).$
@@
Similarly, $t$ centralizes $[wT,T]$, and thus
$\operatorname{Tr}(\operatorname{Ad}_{G}(wt)|_{\operatorname{Lie}([wT,T])})=\operatorname{Tr}(\operatorname{Ad}_{G}(w)|_{\operatorname{Lie}([wT,T])}).$
Since $\mathfrak{t}$ is the direct sum of $\operatorname{Lie}([wT,T])$ and the maximal torus of $H$ (on whith $w$ acts trivially), we may also write this as
$\operatorname{Tr}(\operatorname{Ad}_{G}(w)|_{\operatorname{Lie}([wT,T])}=\operatorname{Tr}(\operatorname{Ad}_{G}(w)|_{\mathfrak{t}})-\operatorname{rank}(H),$
and may then observe that
$\operatorname{Tr}(\operatorname{Ad}_{H}(t[wT,T]))-\operatorname{rank}(H)$
is unchanged if we replace $H$ by its adjoint quotient. ∎
@@
Let $G$ be a compact Lie group with maximal torus $T$, and let $wT\in N_{G}(T)/T$. Then
$\inf_{g\in wT}\operatorname{Tr}\operatorname{Ad}_{G}(g)=\inf_{h\in H}\operatorname{Tr}\operatorname{Ad}_{H}(h)+\operatorname{Tr}(\operatorname{Ad}_{G}(w)|_{\mathfrak{t}})-\operatorname{rank}(H),$
where $H$ is the adjoint group with root system consisting of the $w$-invariant roots of $G$.
@@
If we choose $w\in wT$ to centralize the identity component of $N_{G}([wT,T])/[wT,T]$, then we have
$\inf_{t\in T}\operatorname{Tr}\operatorname{Ad}_{G}(wt)=\inf_{t\in T}\left(\operatorname{Tr}\operatorname{Ad}_{H}(t)+\operatorname{Tr}(\operatorname{Ad}_{G}(w)|_{\mathfrak{t}})-\operatorname{rank}(H)\right);$
pulling the constants out of the infimum and noting that
$\inf_{t\in T}\operatorname{Tr}\operatorname{Ad}_{H}(t)=\inf_{h\in H}\operatorname{Tr}\operatorname{Ad}_{H}(h)$
gives the desired result, once we observe that the roots of $H$ are precisely the $w$-invariant roots of $G$. ∎
In the particular case that $w=c$ corresponds to a diagram automorphism, we find that $H$ is the semisimple group with root system consisting of the $c$-invariant roots of $G$. The action of $c$ permutes the simple factors of $\operatorname{Ad}(G)$, but the roots are contained in single such factors, and thus the root system of $H$ comes only from those components of the Dynkin diagram that are preserved by $c$.
We thus are left with only the indecomposable cases ${}^{2}A_{n}$, ${}^{2}D_{n}$, ${}^{3}D_{4}$, ${}^{2}E_{6}$ to consider. Note that $\operatorname{Ad}(c)|_{\mathfrak{t}}$ is a permutation matrix in the basis corresponding to simple roots, and thus its trace is nothing other than the number of simple roots fixed by $c$.
@@
In the indecomposable cases, the type of $H$ is described in Table 6.
type of $G$ order of $c$ type of $H$ $\operatorname{Tr}(\operatorname{Ad}(c)|_{\mathfrak{t}})-\operatorname{rank}(H)$
@@
Table 6. Subquotient $H$ of $G$ corresponding to a diagram automorphism $c$
@@
We shorten our notation by writing $abcd$ for the root $a\alpha_{1}+b\alpha_{2}+c\alpha_{3}+d\alpha_{4}$. if the automorphism $c$ has order 3, then the only positive roots fixed by $c$ are 0100, $1$, and 1211, which generate a root system of type $A_{2}$. If the automorphism $c$ has order 2, say by interchanging $\alpha_{3}$ and $\alpha_{4}$, then the fixed positive roots are positive linear combinations of 1000, 0100, 0111. Examining the inner products of these roots we find a root system of type $A_{3}=D_{3}$.
@@
For type $E_{6}$ and $c$ of order 2, the fixed simple roots are spanned by
@@
For type $D_{n}$ ($n\geq 5$) and $c$ of order 2, the fixed simple roots are spanned by $\alpha_{1},\ldots,\alpha_{n-2}$ and $\alpha_{n-2}+\alpha_{n-1}+\alpha_{n}$, which span a root system of type $D_{n-1}$ with terminal vertices $\alpha_{1}$, $\alpha_{n-2}$ and $\alpha_{n-2}+\alpha_{n-1}+\alpha_{n}$.
For type $A_{n}$ with $n\geq 2$ and $c$ of order 2, we divide into cases based on whether $n$ is even or odd. If $n$ is even, i.e., $n=2m$ for $m\geq 1$, the fixed positive roots are spanned by the roots
@@
Corollary 13.7 provides the values for $\underline{m}(s)$ and $\overline{m}(s)$ when $s\neq 1$, relying on the $s=1$ case of Theorem 2.3 for the group $H$, as we now calculate.
For example, when $G$ has type $E_{6}$ and $c$ has order 2, $\operatorname{Tr}\operatorname{Ad}$ on $D_{4}$ ranges from $-4$ to 28. Since $\operatorname{Tr}(\operatorname{Ad}(c)|_{\mathfrak{t}})-\operatorname{rank}(H)=-2$, we find that the $\operatorname{Tr}\operatorname{Ad}$ on $cG$ varies from $-6$ to 26.
When $G$ has type $D_{4}$ and $c$ has order 3, $\operatorname{Tr}\operatorname{Ad}$ on $A_{2}$ ranges from $-1$ to $8$, so $\operatorname{Tr}\operatorname{Ad}$ on $cG$ varies from $-2$ to 7.
When $G$ has type $D_{n}$ and $c$ has order 2, $\operatorname{Tr}\operatorname{Ad}$ on $D_{n-1}$ has maximum $32n^{2}-5n+$ 3, so $\operatorname{Tr}\operatorname{Ad}$ on $cG$ has maximum $22n^{2}-5n+$ 2. The minimum of $\operatorname{Tr}\operatorname{Ad}$ on $D_{n-1}$ is $-(n-1)$ if $n$ is odd and $2-(n-1)$ if $n$ is even, whence the minimum of $\operatorname{Tr}\operatorname{Ad}$ on $cG$ is $-n$ if $n$ is odd and $2-n$ if $n$ is even.
When $G$ has type $A_{n}$ and $c$ has order 2, $\operatorname{Tr}\operatorname{Ad}$ on $A_{1}^{\times m}$ ranges from $-m$ to $3m$. If $n=2m$, then the correction term is $-m$, so $\operatorname{Tr}\operatorname{Ad}$ on $cG$ ranges from $-2m=-n$ to $2m=n$. If $n=2m-1$, then the correction term is $1-m$, so $\operatorname{Tr}\operatorname{Ad}$ on $cG$ ranges from $-2m+1=-n$ to $22m+1=n+$ 2. ∎
@@
Let the random element $w\in G$ be chosen uniformly from a coset of $T$ in $N_{G}(T)$. Then $\operatorname{Tr}(\operatorname{Ad}(w))$ has mean $\operatorname{Tr}(\operatorname{Ad}(w)|_{\mathfrak{t}})$ and variance the number of roots fixed by $w$.

### v4

@@ 11. Type $C$
Note that the eigenvalues of an element of $\operatorname{Sp}_{2n}$ coming from the compact form have $\bar{s}_{j}=1/s_{j}$, so each $t_{j}$ is real. It follows that the power sums $p_{d}=\sum_{i}t_{i}^{d}$ have $p_{1}^{2}$ and $p_{2}$ positive. The expression $\sum t_{j}^{2}+\sum t_{j}t_{k}$ is the degree 2 complete homogeneous symmetric function in the $t_{j}$’s, so it is $(p_{1}^{2}+p_{2})/2\geq 0$. Then (11.1) gives the required bound $\operatorname{Tr}\operatorname{Ad}(g)\geq-(\operatorname{rank}G)$, proving Theorem 2.3 for type $C$.
@@
The claim for type $C_{n}$ in Proposition 2.5 now follows directly from Lemma 10.2.
@@ 12. Type $B$
The highest short root representation $\rho$ of $\operatorname{SO}(2n+1)$ is the natural representation $\mathbb{R}^{2n+1}$, so
@@ 13. Components other than the identity
To complete the proof of Theorem 2.3, it remains to compute $\underline{m}(s)$ and $\overline{m}(s)$ when $s\neq 1$. In the following results, we consider a compact Lie group $G$, and our interest is in the case where $G$ is not connected.
A subquotient $H$
Let $T$ be a maximal torus in $G$. Since $T$ is abelian, for $w\in N_{G}(T)$ the function $T\to T$ defined by $t\mapsto[w,t]:=(wtw^{-1})t^{-1}$ is a homomorphism. (This same notion appeared in the proof of Lemma 7.2.) We denote its image, a closed subgroup of $T$, by $[w,T]$.
@@
Let $w\in N_{G}(T)$, and let $\rho:G\to\operatorname{GL}(V)$ be a representation of $G$. Then for all $t\in T$,
$\operatorname{Tr}(\rho(wt))=\operatorname{Tr}(\rho(wt)|_{V^{[w,T]}}).$
@@
Over $\mathbb{C}$, $V$ is a direct sum of weight spaces $V_{\lambda}$ for weights $\lambda$. If $w\cdot\lambda\neq\lambda$, then the action of $w$ permutes the weight spaces for weights in the orbit of $\lambda$, so those summands together contribute 0 to the trace. We thus find that
$\operatorname{Tr}(\rho(w))=\sum_{\lambda:w\cdot\lambda=\lambda}\operatorname{Tr}(\rho(w)|_{V_{\lambda}})=\operatorname{Tr}(\rho(w)|_{V^{\prime}}),$
where $V^{\prime}$ is the direct sum of the weight spaces corresponding to $w$-invariant weights. A weight is $w$-invariant if and only if $\lambda(wtw^{-1})=\lambda(t)$ for all $t\in T$, if and only if $\lambda(wtw^{-1}t^{-1})=1$ for all $t\in T$, if and only if $\lambda|_{[w,T]}=1$. In other words, $V^{\prime}$ is nothing other than the subspace of $V$ of vectors fixed by $[w,T]$, as required. ∎
In particular, when optimizing a character over a coset $wT$, we may restrict to the subspace $V^{[w,T]}$. This is naturally a representation of a subquotient of $G$, namely $N_{G}([w,T])/[w,T]$. Call the identity component of this group $H$; equivalently, $H=N_{G}([w,T])^{\circ}/[w,T]$. (This $H$ will be used in this section and reappear in §15 as part of the proof of Cor. 2.4.)
Certainly $T$ normalizes its subtorus $[w,T]$. For $t\in T$ we have $w[w,t]w^{-1}=[w,wtw^{-1}]\in[w,T]$, so $wT$ is contained in $N_{G}([w,T])$ and every element of $wT$ normalizes $H$.
The Lie algebra of $N_{G}([w,T])$ is the $[w,T]$-invariant subspace of $\mathfrak{g}$. Its roots (i.e., the nonzero weights of $\mathfrak{g}^{[w,T]}$) are the $w$-invariant roots of $G$, while its maximal torus is $T/[w,T]=H_{0}(\langle w\rangle;T)$.
@@
There is an element $w^{\prime}\in wT$ that centralizes $H$.
@@
Write $\operatorname{Ad}$ for the homomorphisms
$H\to\operatorname{Aut}(H)\leftarrow N_{G}([w,T])$
that send an element to conjugation by that element. Since $\operatorname{Ad}(w)\in\operatorname{Aut}(H)$ fixes the maximal torus $T/[w,T]$ of $H$ elementwise, $\operatorname{Ad}(w)$ equals conjugation by an element of $T/[w,T]$ by the following lemma. Since $\operatorname{Ad}(T)=\operatorname{Ad}(T/[w,T])$, we have $\operatorname{Ad}(w)=\operatorname{Ad}(t)$ for some $t\in T$, and $w^{\prime}:=wt^{-1}$ centralizes $H$ as required. ∎
Here is the promised lemma, which is implicit in standard references such as [Bou05, §IX.4.10].
Lemma 13.3.
Suppose the compact Lie group $G$ is connected. If an automorphism $\phi$ of $G$ fixes the maximal torus $T$ elementwise, then there is a $t\in T$ such that $\phi(g)=tgt^{-1}$ for $g\in G$.
Proof.
Since $\phi$ normalizes $T$, it acts on $\operatorname{Hom}(T,S^{1})$, and this action is trivial since $\phi$ fixes $T$ elementwise. Therefore, the automorphism of the Dynkin diagram of $G$ induced by $\phi$ is trivial and it follows from [Bou05, §IX.4.10, Eq. (16)] that there is some $x\in G$ such that $\phi(g)=xgx^{-1}$ for all $g\in G$. Since $x$ centralizes $T$, it belongs to $T$. ∎
We now return to the setting where $G$ need not be connected. When $w$ centralizes $H$, Lemma 13.1 has a particularly nice consequence.
@@
$\operatorname{Tr}(\operatorname{Ad}_{G}(wt))=\operatorname{Tr}(\operatorname{Ad}_{H}(t))+\operatorname{Tr}(\operatorname{Ad}_{G}(w)|_{\mathfrak{t}})-\operatorname{rank}(H)$
for $t\in T$.
@@
Let $H^{+}$ be the normalizer of $[w,T]$, so that its Lie algebra is $\mathfrak{g}^{[w,T]}$. Then by Lemma 13.1
@@
$\operatorname{Tr}(\operatorname{Ad}_{H^{+}}(wt))$ $=\operatorname{Tr}(\operatorname{Ad}_{H}(wt))+\operatorname{Tr}(\operatorname{Ad}_{H^{+}}(wt)|_{\operatorname{Lie}([w,T])})$
$=\operatorname{Tr}(\operatorname{Ad}_{H}(wt))+\operatorname{Tr}(\operatorname{Ad}_{G}(wt)|_{\operatorname{Lie}([w,T])}).$
@@
Similarly, $t$ centralizes $[w,T]$, so
$\operatorname{Tr}(\operatorname{Ad}_{G}(wt)|_{\operatorname{Lie}([w,T])})=\operatorname{Tr}(\operatorname{Ad}_{G}(w)|_{\operatorname{Lie}([w,T])}).$
Since $\mathfrak{t}$ is the direct sum of $\operatorname{Lie}([w,T])$ and the maximal torus of $H$ (on which $w$ acts trivially), we may also write this as
$\operatorname{Tr}(\operatorname{Ad}_{G}(w)|_{\operatorname{Lie}([w,T])})=\operatorname{Tr}(\operatorname{Ad}_{G}(w)|_{\mathfrak{t}})-\operatorname{rank}(H).\qed$
@@
Let $G$, $T$, $wT$, and $H$ be as above. Then
$\inf_{g\in wT}\operatorname{Tr}\operatorname{Ad}_{G}(g)=\left(\inf_{h\in H}\operatorname{Tr}\operatorname{Ad}_{H}(h)\right)+\operatorname{Tr}(\operatorname{Ad}_{G}(w)|_{\mathfrak{t}})-\operatorname{rank}(H).$
The same statement holds with $\inf$ replaced by $\sup$.
@@
We may replace $w$ with $wt$ for $t\in T$, and so (by Lemma 13.2) assume that $w$ centralizes $H$. Then we have by Lemma 13.4:
$\inf_{t\in T}\operatorname{Tr}\operatorname{Ad}_{G}(wt)=\inf_{t\in T}\left(\operatorname{Tr}\operatorname{Ad}_{H}(t)+\operatorname{Tr}(\operatorname{Ad}_{G}(w)|_{\mathfrak{t}})-\operatorname{rank}(H)\right).$
Since $H$ is compact every $h\in H$ is conjugate to an element in the image of $T$, so
$\inf_{t\in T}\operatorname{Tr}\operatorname{Ad}_{H}(t)=\inf_{h\in H}\operatorname{Tr}\operatorname{Ad}_{H}(h),$
which gives the desired result for $\inf$. The same argument works verbatim with $\inf$ replaced by $\sup$. ∎
Consider now the case where $G$ is simple and $w$ acts via a diagram automorphism, relative to some pinning of $G^{\circ}$ with respect to the maximal torus $T$. Then $\operatorname{Ad}(w)|_{\mathfrak{t}}$ is a permutation matrix in the basis corresponding to simple roots, and thus its trace equals the number of simple roots fixed by $w$. To effectively apply Cor. 13.5 in this case, we need to describe the subquotient $H$, which we now do.
@@
Let $G$ and $w$ be as in the preceding paragraph. If $w\neq 1$, then the type of $H$ is described in Table 6.
type of $G$ order of $c$ type of $H$ $\operatorname{Tr}(\operatorname{Ad}(w)|_{\mathfrak{t}})-\operatorname{rank}(H)$
@@
Table 6. Subquotient $H$ of $G$ corresponding to a diagram automorphism $w\neq 1$
@@
We shorten our notation by writing $abcd$ for the root $a\alpha_{1}+b\alpha_{2}+c\alpha_{3}+d\alpha_{4}$. if the automorphism $w$ has order 3, then the only positive roots fixed by $w$ are 0100, $1$, and 1211, which generate a root system of type $A_{2}$. If the automorphism $w$ has order 2, say by interchanging $\alpha_{3}$ and $\alpha_{4}$, then the fixed positive roots are positive linear combinations of 1000, 0100, 0111. Examining the inner products of these roots we find a root system of type $A_{3}=D_{3}$.
@@
For type $E_{6}$ and $w$ of order 2, the fixed simple roots are spanned by
@@
For type $D_{n}$ ($n\geq 5$) and $w$ of order 2, the fixed simple roots are spanned by $\alpha_{1},\ldots,\alpha_{n-2}$ and $\alpha_{n-2}+\alpha_{n-1}+\alpha_{n}$, which span a root system of type $D_{n-1}$ with terminal vertices $\alpha_{1}$, $\alpha_{n-2}$ and $\alpha_{n-2}+\alpha_{n-1}+\alpha_{n}$.
For type $A_{n}$ with $n\geq 2$ and $w$ of order 2, we divide into cases based on whether $n$ is even or odd. If $n$ is even, i.e., $n=2m$ for $m\geq 1$, the fixed positive roots are spanned by the roots
@@
Let $G$ be a compact Lie group that is simple and adjoint. Pick a maximal torus $T$ of $G$ and a pinning of $G$ with respect to $T$. The pinning identifies $\operatorname{Aut}(G)$ with $G\rtimes A$ for $A$ the group of automorphisms of the Dynkin diagram of $G$ [Bou05, §IX.4.10, Cor.], and in particular $A$ is identified with $\operatorname{Out}(G)$. For $s\neq 1$ in $\operatorname{Out}(G)$, we have a corresponding element $w\in A$. To prove Theorem 2.3, our task is to find the extreme values of $\operatorname{Tr}\operatorname{Ad}(gw)$ for $g\in G$.
A result of de Siebenthal [dS56, Th. on p. 57] says that the connected component $Gs$ of $\operatorname{Aut}(G)$ is a union of $G$-conjugates of $Tw$, so we are reduced to finding the extreme values of $\operatorname{Tr}\operatorname{Ad}(wt)$ for $t\in T$. For this, we may apply Cor. 13.5, where the extreme values of $\operatorname{Ad}_{H}(h)$ are provided by the $s=1$ case of Theorem 2.3 (which we have already proved) and $\operatorname{Tr}(\operatorname{Ad}_{G}(w)|_{\mathfrak{t}})-\operatorname{rank}(H)$ is provided by Table 6.
For example, when $G$ has type $E_{6}$ and $w$ has order 2, $\operatorname{Tr}\operatorname{Ad}$ on $D_{4}$ ranges from $-4$ to 28. Since $\operatorname{Tr}(\operatorname{Ad}(w)|_{\mathfrak{t}})-\operatorname{rank}(H)=-2$, we find that the $\operatorname{Tr}\operatorname{Ad}$ on $Gw$ varies from $-6$ to 26.
When $G$ has type $D_{4}$ and $w$ has order 3, $\operatorname{Tr}\operatorname{Ad}$ on $A_{2}$ ranges from $-1$ to $8$, so $\operatorname{Tr}\operatorname{Ad}$ on $Gw$ varies from $-2$ to 7.
When $G$ has type $D_{n}$ and $w$ has order 2, $\operatorname{Tr}\operatorname{Ad}$ on $D_{n-1}$ has maximum $32n^{2}-5n+$ 3, so $\operatorname{Tr}\operatorname{Ad}$ on $Gw$ has maximum $22n^{2}-5n+$ 2. The minimum of $\operatorname{Tr}\operatorname{Ad}$ on $D_{n-1}$ is $-(n-1)$ if $n$ is odd and $2-(n-1)$ if $n$ is even, whence the minimum of $\operatorname{Tr}\operatorname{Ad}$ on $Gw$ is $-n$ if $n$ is odd and $2-n$ if $n$ is even.
When $G$ has type $A_{n}$ and $w$ has order 2, $\operatorname{Tr}\operatorname{Ad}$ on $A_{1}^{\times m}$ ranges from $-m$ to $3m$. If $n=2m$, then the correction term is $-m$, so $\operatorname{Tr}\operatorname{Ad}$ on $Gw$ ranges from $-2m=-n$ to $2m=n$. If $n=2m-1$, then the correction term is $1-m$, so $\operatorname{Tr}\operatorname{Ad}$ on $Gw$ ranges from $-2m+1=-n$ to $22m+1=n+$ 2. ∎
@@
Let the random element $w^{\prime}\in G$ be chosen uniformly from a coset $wT$ of $T$ in $N_{G}(T)$. Then $\operatorname{Tr}(\operatorname{Ad}(w^{\prime}))$ has mean $\operatorname{Tr}(\operatorname{Ad}(w)|_{\mathfrak{t}})$ and variance the number of roots fixed by $w$.